A piano tuner using a 264 Hz tuning fork hears 6 beats per second while playing the tuning fork and a piano key at the same time. If the tuner switches to a 268 Hz tuning fork, what are the possible beat frequencies that could be detected?
Answer #1

The frequency of the piano key may be (264 - 6) = 258 Hz or (264 + 6) = 270 Hz

Therefore, with 268 Hz tuning form, the possible beat frequencies are

(268 - 258) = 10 Hz. Answer.

Or

(270 - 268) = 2 Hz. Answer